OpenStudy (anonymous):

Solve the following word problem by using an equation and factoring. The sum of a positive number and its square is 20. What is the number? A. 1 B. 2 C. 3 D. 4 E. 5

OpenStudy (anonymous):

if you call the number \(x\) then you are trying to solve \(x^2+x=20\)

OpenStudy (anonymous):

you can start with \(x^2+x-20=0\) and then factor as suggested do you know how to factor this one?

OpenStudy (anonymous):

@satellite73 not really ;/

OpenStudy (anonymous):

come up with two numbers whose product is \(-20\) and that add to \(1\) one will be positive, the other negative

OpenStudy (anonymous):

if you don't get it let me know

OpenStudy (anonymous):

so is the answer E?

OpenStudy (anonymous):

i meant can you factor the answer is not E

OpenStudy (anonymous):

I still don't get it.

OpenStudy (anonymous):

you can factor as \((x+5)(x-4)=0\)

OpenStudy (anonymous):

then either \(x+5=0\) giving \(x=-5\) or \(x-4=0\) so \(x=4\)

OpenStudy (anonymous):

since you are looking only for the positive answer, use \(x=4\)

OpenStudy (anonymous):

So the answer is D because it's 4?

OpenStudy (anonymous):

and we can check that it works \[4^2=4=16+4=20\]

OpenStudy (anonymous):

yes, it is D